# AuToMATo: An Out-Of-The-Box Persistence-Based Clustering Algorithm

### v2

@@ 4.4 Results and interpretation
Table LABEL:table:summary_of_results shows the average Fowlkes-Mallows score of each algorithm across all benchmarking data sets. Moreover, to account for the randomized component of AuToMATo, this table reports the mean and the standard deviation of the Fowlkes-Mallows score across the ten runs. For those benchmarking data sets that come with more than one ground truth, we included only the best score of the respective algorithm. Similarly, we included only the best performing parameter selection for those algorithms that we ran with varying distance thresholds (which, of course, skews the comparison in favor of those algorithms). As Table LABEL:table:summary_of_results shows, AuToMATo outperforms each clustering algorithm on average across all data sets, thus showing that it is indeed a versatile and powerful out-of-the-box clustering algorithm. In particular, AuToMATo outperforms the TTK-algorithm, which also build on ToMATo.
@@
Of course, we do not claim that AuToMATo outperforms each of the algorithms above on every single data set. Indeed, seeing that we ran DBSCAN with a wide range of choices for the distance threshold parameter, it is to be expected that the best choice of parameter should often outperform AuToMATo. However, on most data sets where this is the case, the results from AuToMATo are still competitive. In fact, there are a significant number of instances where AuToMATo outperforms DBSCAN for all parameter selections, in some cases by a lot. This is illustrated in Figure 4, where we show the Fowlkes-Mallows score of AuToMATo and each of the best and worst performing parameter choice of DBSCAN across all data sets.
Figure 4: Comparison of AuToMATo and DBSCAN.
The scores of our experiments are reported in Tables LABEL:table:aut_vs_dbscan through LABEL:table:aut_vs_ttk in Appendix A.2, and each of these tables is accompanied by a figure similar to Figure 4.
@@ 5 Applications of AuToMATo in Combination with Mapper
The goal of Mapper [SMC+07] is to approximate the Reeb graph of a manifold $M$ based on a sample from $M$. The input is a point cloud $P$ with a filter function $P\rightarrow\mathbb{R}$, a collection of overlapping intervals $\mathcal{U}=\{U_{1},\ldots,U_{n}\}$ covering $\mathbb{R}$ and a clustering algorithm. For each $U_{i}\in\mathcal{U}$, Mapper runs the clustering algorithm on the data points in the preimage $f^{-1}(U_{i})$, creating a vertex for each cluster. Two vertices are then connected by an edge if the corresponding clusters (in different preimages) have some data points in common, yielding a graph that represents the shape of the data set.
We ran the Mapper implementation of giotto-tda [TLT+20] on a synthetic two-dimensional data set consisting of noisy samples from two concentric circles (see Figure 5(a)) with projection onto the $x$-axis as the filter function. We ran Mapper on the same interval cover with three different choices of clustering algorithms: AuToMATo, DBSCAN, and HDBSCAN (the latter two with their respective default parameters). As can be seen in Figure 5(b), using DBSCAN, we get many unwanted edges in the graph. HDBSCAN performs better, giving two cycles with some extra loops. The output of Mapper with AuToMATo is exactly the Reeb graph of two circles.
@@
We further tested the combination of Mapper with AuToMATo on one of the standard applications of Mapper: the Miller-Reaven diabetes data set, where Mapper can be used detect two strains of diabetes that correspond to “flares” in the data set (see [SMC+07, Section 5.1] for details). [^8] The data set is available as part of the “locfit” R-package [Loa24]. As can be seen in Figure 6, AuToMATo performs well in this task; the graphs show a central core of vertices corresponding to healthy patients, and two flares corresponding to the two strains of diabetes. We were not able to reproduce this using DBSCAN or HDBSCAN; Figure 6 shows the output of Mapper with these algorithms with their respective default parameters.
@@ 6 Discussion
We briefly outline some limitations of AuToMATo. As mentioned in Remark 3.1, AuToMATo comes with a choice of default values for its parameters that make it perform well in experiments. In particular, AuToMATo resorts to the default values as implemented in ToMATo for the choice of neighborhood graph and density estimators. In ToMATo, the options for the neighborhood graph estimators are the $\delta$-Rips graph and the $k$-nearest neighbor graph, relying on parameters $\delta>0$ and $k\geq 1$, respectively; the options for the density estimators are the kernel density estimator and the distance-to-measure density estimator, which in turn rely parameters $h>0$ and $m\in(0,1)$, respectively (see the footnotes in Section 2.2 for their definitions). While for most data sets that we ran our experiments on these default estimators yielded good results, there are cases where adjusting them before running AuToMATo improves the clustering performance. Nevertheless, since finding a priori optimal parameters for the neighborhood graph and density estimators for a given data set is largely an open problem, we chose to stick to the default values for these estimators. Indeed, there are heuristics for the selection of the bandwidth in kernel density estimation (see [CSTL21, Section 4.1]) and the smoothing parameter in distance-to-measure density estimation (see [CFL+17, Section 7.1]), but these methods either lack theoretical justification, or would require running AuToMATo multiple times on the same data set with different parameters and selecting the best ones a posteriori, effectively rendering AuToMATo a non-parameter-free clustering algorithm.
Optimizing the choice of the neighborhood graph and density estimators is an aspect of AuToMATo that we plan to pursue in future work. Moreover, we plan to improve the currently experimental feature for outlier creation in AuToMATo discussed at the end of Section 3.2. Finally, it is natural to ask whether the results from [CMO18] on optimal parameter selection in the Mapper algorithm can be adapted to the scenario where Mapper uses AuToMATo as its clustering algorithm.
@@ Appendix A Appendix
A.1 Complexity analysis of Algorithm 1
Recall from [CGOS13, Section 2] that, if an estimated density and a neighborhood graph are provided, ToMATo has a worst-case time complexity in $O(n\log(n)+m\alpha(n))$, where $n$ and $m$ are the number of vertices and edges of the neighborhood graph, respectively, and $\alpha$ denotes the inverse Ackermann function (note that $n$ equals the number of data points). By default, ToMATo (and hence AuToMATo) works with the $k$-nearest neighbor graph and distance-to-measure density estimators, where the latter relies itself on the $k$-nearest neighbor graph (each with $k=10$). Taking into account the known complexity bound $O(nd)$ for the creation of the $k$-nearest neighbor graph (where $d$ is the dimensionality of the data), and using the fact that $m\in O(n)$ for this graph, this leads to a worst-case time complexity in $O(nd+n\log(n))$ for a single run of ToMATo. Creating the bootstrap samples $X_{i}^{*}$, $i=1,\dots,B$, has complexity in $O(Bn)$; computing the values $\sqrt{n}W_{\infty}^{\mathrm{fin}}(\mathcal{D}_{i}^{*},\mathcal{D})$, $i=1,\dots,B$, has worst-case complexity $O(BN^{1.5}\log(N))$ (see e.g. [EIK01]; here $N$ denotes the maximal number of off-diagonal points across all relevant persistence diagram), and sorting them has worst case complexity in $O(B\log(B))$. Combined, this leads to a worst-case complexity for AuToMATo in $O(B(nd+n\log(n)+N^{1.5}\log(N))+B\log(B))$. Using that $B$ is a constant, we obtain the runtime claimed in the main body.
@@ A.2 Benchmarking results
In this subsection we report the Fowlkes-Mallows scores coming from comparing AuToMATo to the other clustering algorithms, as explained in Section 4. For those benchmarking data sets that come with more than one ground truth, we report the scores for each of those, and different ground truths are indicated by the last digit in the data set name. Moreover, each table is sorted according to increasing difference in clustering performance of AuToMATo and the respective clustering algorithm that AuToMATo is being compared against. As is customary, we indicate the score stemming from the best performing clustering algorithm in bold.

### v3

@@ 4.4 Results and interpretation
Table LABEL:table:summary_of_results shows the average Fowlkes-Mallows score of each algorithm across all benchmarking data sets; for AuToMATo, it shows the average and the standard deviation across the ten runs. For those benchmarking data sets that come with more than one ground truth, we included only the best score of the respective algorithm. Similarly, we included only the best performing parameter selection for those algorithms that we ran with varying distance thresholds (which, of course, skews the comparison in favor of those algorithms). As Table LABEL:table:summary_of_results shows, AuToMATo outperforms each clustering algorithm on average across all data sets, thus showing that it is indeed a versatile and powerful out-of-the-box clustering algorithm. In particular, AuToMATo outperforms the TTK-algorithm, which also build on ToMATo.
@@
The scores of our experiments are reported in Tables LABEL:table:aut_vs_dbscan through LABEL:table:aut_vs_ttk in Appendix A.2. As an illustration, Figure 3 shows that the best choice of parameter for DBSCAN sometimes outperforms AuToMATo, which is to be expected. However, on most data sets where this is the case, the results from AuToMATo are still competitive, and there is a significant number of instances where AuToMATo outperforms DBSCAN for all parameter selections, in some cases by a lot.
Figure 3: Fowlkes-Mallows score of AuToMATo and DBSCAN across benchmarking data sets. The shading of “automato_mean” indicates the standard deviation of the score across the ten runs.
@@ 5 Applications of AuToMATo in combination with Mapper
The goal of Mapper (Singh et al., 2007) is to approximate the Reeb graph of a manifold $M$ based on a sample from $M$. The input is a point cloud $P$ with a filter function $P\rightarrow\mathbb{R}$; a collection of overlapping intervals $\mathcal{U}=\{U_{1},\ldots,U_{n}\}$ covering $\mathbb{R}$; and a clustering algorithm. For each $U_{i}\in\mathcal{U}$, Mapper runs the clustering algorithm on the data points in the preimage $f^{-1}(U_{i})$, creating a vertex for each cluster. Two vertices are then connected by an edge if the corresponding clusters (in different preimages) have some data points in common, yielding a graph that represents the shape of the data set.
We ran the Mapper implementation of giotto-tda (Tauzin et al., 2020) on a synthetic two-dimensional data set consisting of noisy samples from two concentric circles (see Figure 4(a)) with projection onto the $x$-axis as the filter function. We ran Mapper on the same interval cover with three different choices of clustering algorithms: AuToMATo, DBSCAN, and HDBSCAN. As can be seen in Figure 4(b), using DBSCAN, we see many unwanted edges in the graph. HDBSCAN performs better, giving two cycles with some extra loops. The output of Mapper with AuToMATo is exactly the Reeb graph of two circles.
@@
We further tested the combination of Mapper with AuToMATo on one of the standard applications of Mapper: the Miller-Reaven diabetes data set, where Mapper can be used detect two strains of diabetes that correspond to “flares” in the data set (see Singh et al. (2007, Section 5.1) for details). [^9] The data set is available as part of the “locfit” R-package (Loader, 2024). As can be seen in Figure 5, AuToMATo performs well in this task; the graphs show a central core of vertices corresponding to healthy patients, and two flares corresponding to the two strains of diabetes. We were not able to reproduce this using DBSCAN or HDBSCAN; Figure 5 shows the output of Mapper with these algorithms with their respective default parameters.
@@ 6 Discussion
We briefly outline some limitations of AuToMATo. AuToMATo comes with a choice of default values for its parameters.
Optimizing the choice of the neighborhood graph and density estimators is an aspect of AuToMATo that we plan to pursue in future work. Moreover, we plan to improve the currently experimental feature for outlier creation in AuToMATo discussed at the end of Section 3.2. Finally, it is natural to ask whether the results from Carrière et al. (2018) on optimal parameter selection in the Mapper algorithm can be adapted to the scenario where Mapper uses AuToMATo as its clustering algorithm.
Acknowledgments
The first author was supported by the Swiss National Science Foundation (project no. 209413).
@@ Appendix A Appendix
A.1 About the choice of data sets
As explained in Section 4.2, we chose to include the data set named windows from the battery named wut, but not the data set named olympic from the same battery. Those are illustrated in Figure 6. In that figure, the data points are colored according to the ground truth clustering.
@@ A.2 Benchmarking results
In this subsection we report the Fowlkes-Mallows scores coming from comparing AuToMATo to the other clustering algorithms, as explained in Section 4. For those benchmarking data sets that come with more than one ground truth, we report the scores for each of those, and different ground truths are indicated by the last digit in the data set name. Moreover, each table is sorted according to increasing difference in clustering performance of AuToMATo and the respective clustering algorithm that AuToMATo is being compared against. As is customary, we indicate the score stemming from the best performing clustering algorithm in bold. Finally, each of the table is accompanied by a graph similar to the one depicted in Figure 3. Note that, in particular, that those figures indicate only the score corresponding to the ground truth on which the respective clustering algorithm performs best on.